### v2

# DyBBT: Dynamic Balance via Bandit inspired Targeting for Dialog Policy with Cognitive Dual Systems

Abstract
Task oriented dialog systems often rely on static exploration strategies that do not adapt to dynamic dialog contexts, leading to inefficient exploration and suboptimal performance. We propose DyBBT, a novel dialog policy learning framework that formalizes the exploration challenge through a structured cognitive state space $\mathcal{C}$ that captures dialog progression, user uncertainty, and slot dependency. DyBBT proposes a bandit inspired meta-controller that dynamically switches between a fast intuitive inference (System 1) and a slow deliberative reasoner (System 2) based on real-time cognitive states and visitation counts. Extensive experiments on single- and multi-domain benchmarks show that DyBBT achieves SOTA performance in success rate, efficiency, and generalization, with human evaluations confirming that its decisions are well aligned with expert judgment. The code is available at Anonymous Github.
DyBBT: Dynamic Balance via Bandit inspired Targeting for Dialog Policy with Cognitive Dual Systems
Shuyu Zhang1thanks: These authors contributed equally to this work.,  Yifan Wei2 *,  Jialuo Yuan3,  Xinru Wang4,  Yanmin Zhu1thanks: Corresponding Author.,   Bin Li5 †  Yujie Liu6, 1Department of Computer Science and Engineering, Shanghai Jiao Tong University 2School of Computer Science and Engineering, Beihang University 3Stanford University 4Data Analytics for Business, University of Sydney 5Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences 6School of Information Engineering, Beijing Institute of Graphic Communication carsonz@sjtu.edu.cn, weiyifan@buaa.edu.cn, xinruwang1999@outlook.com yzhu@cs.sjtu.edu.cn, yjl705@stanford.edu, b.li2@siat.ac.cn 18764008899@163.com
1 Introduction
“The affordances of the environment are what it offers the animal, what it provides or furnishes, for good or ill.”
— The Ecological Approach to Visual Perception (Gibson, 1979)
Figure 1: A comparison of exploration strategies for dialog policy learning. Traditional DRL relies on static heuristics incapable of adapting to dynamic dialog contexts. EIERL uses population based evolution but struggles in complex tasks. DyBBT solves the adaptive exploration challenge by cognitive meta-controller to achieve a principled balance between efficiency and robustness.
Task-oriented dialog systems (TODS) assist users in achieving specific goals through multi-turn interactions. Dialog policy learning is typically formulated as a sequential decision making problem addressed with Deep Reinforcement Learning (DRL) (Nachum et al., 2017; Silver et al., 2014). However, it is fundamentally bottlenecked by the exploration exploitation dilemma: balancing the exploitation of known rewards against the exploration of unknown actions to discover better strategies. In TODS, this dilemma is exacerbated by the dynamic and partially observable context, characterized by quantifiable cognitive features such as dialog progress, user intent entropy, and slot dependencies (Peng et al., 2017; Wen et al., 2017). The features govern the cost benefit of exploration: early in a dialog, high entropy makes information gathering valuable, and later high slot dependency makes exploitation critical to avoid constraint violations (Qin et al., 2023; Zhao et al., 2024).
Existing methods for enhancing exploration in TODS remain misaligned with this dynamic cognitive reality. As illustrated in Fig. 1, traditional DRL methods rely on static heuristics like $\epsilon$-greedy (Niu et al., 2024), which cannot adapt to shifting exploration needs across dialog phases. Evolutionary methods like EIERL (Zhao et al., 2025) enable global search via population based optimization but struggle to scale in complex multi-domain scenarios. LLM-based policies (Zhang et al., 2024; He et al., 2022) or reasoning techniques like Tree of Thoughts (Yao et al., 2023) support deep deliberative planning but incur prohibitive computational overhead and lack a principled mechanism to trigger such costly reasoning only when necessary. This misalignment reveals a key research question: How can we design a dialog policy that dynamically perceives cognitive affordances to balance exploration and exploitation?
Inspired by Gibson’s affordances theory, we propose that the dialog environment presents a dynamic landscape of exploration opportunities which an effective policy must perceive and act upon. We introduce DyBBT, a novel framework that grounds decisions in an interpretable cognitive state space $\mathcal{C}$ that captures dialog progress, user uncertainty, and slot dependency. It employs a lightweight meta-controller that dynamically switches between a fast System 1 for routine decisions and a slow System 2 for costly deliberation, based on real-time cognitive signals. This design ensures expensive reasoning is invoked only when the cognitive state signals high epistemic uncertainty or low confidence, addressing the core limitations of previous methods.
In summary, our contributions are: (1) Formalizing the TODS exploration challenge via a structured cognitive state space $\mathcal{C}$ (Sec. 3.1); (2) Proposing DyBBT, a novel framework with a meta-controller to dynamically balance between fast and slow reasoning (Sec. 3.1.2); (3) Demonstrating SOTA performance and human aligned decisions through extensive experiments (Sec. 4).
2 Related Work
2.1 DRL for Dialog Policy Learning
Deep Reinforcement Learning (DRL) has become a dominant paradigm for dialog policy optimization due to its capacity for sequential decision making. Early work applied value based methods (Peng et al., 2018) and policy gradient algorithms (Silver et al., 2014) to TODS, with Proximal Policy Optimization (PPO) (Schulman et al., 2017) later adopted for improved stability. A key limitation of these methods is their reliance on static exploration strategies, such as $\epsilon$-greedy or entropy bonus, which cannot adapt to the dynamic uncertainty and structural complexity of multi-domain dialogs (Kwan et al., 2023; Jia et al., 2024). Recent efforts have incorporated Bayesian reasoning (Lee et al., 2023), meta-learning (Li et al., 2024; Liang et al., 2024), and cascading RL (Du et al., 2024) to enable more adaptive exploration. While promising, these approaches often lack an explicit and interpretable representation of the internal dialog state that directly governs exploration, the gap that our structured cognitive state space $\mathcal{C}$ aims to fill.
2.2 Evolutionary Exploration Methods
To overcome static exploration, research has diverged into population based optimization and principled exploration theory. Population based methods like Evolutionary RL (EIERL)  (Zhao et al., 2025) enhance diversity but scale poorly with dialog complexity  (Sigaud, 2023) and lack real time adaptation  (Lin et al., 2025). Theoretically, bandit algorithms (UCB  (Garivier and Moulines, 2011), contextual  (Foster and Rakhlin, 2020), hierarchical RL  (Rohmatillah and Chien, 2023)) and posterior sampling (PSRL  (Chen et al., 2020)) formalize exploration. However, their direct application to dialog POMDPs is hindered by non-stationarity and high dimensionality. Both lines of work lack a mechanism to perceive and respond to the dynamic cognitive affordances such as shifting uncertainty within a dialog. DyBBT bridges this gap by pragmatically adapting bandit inspiration to a learned, low dimensional cognitive state space $\mathcal{C}$, offering a tractable bridge between classical theory and sequential dialog complexity.
2.3 Dual System Reasoning Architectures
Inspired by dual process theory (Krämer, 2014), recent work combines fast processing (System 1) with slow reasoning (System 2) for mathematical reasoning (Shi et al., 2024) and common sense inference (Yu et al., 2025). In TODS, large language models (LLMs) serve as powerful function approximators (Yi et al., 2025), acting as intuitive generators (Ying et al., 2024) and deliberative reasoners (Ma et al., 2025). Frameworks like the Dynamic Dual Process Transformer (He et al., 2024) explicitly model this interaction for policy learning. However, existing switching mechanisms often rely on static heuristics, such as fixed turn counts (Qin et al., 2023) or pre-defined confidence thresholds (Yao et al., 2023), which lack adaptability and theoretical grounding in exploration. DyBBT addresses this by introducing a meta-controller over a cognitive state space, dynamically triggering System 2 based on visitation counts and parametric uncertainty, offering a principled and efficient alternative to heuristic switching.
3 Methodology
DyBBT formulates dialog exploration as a tractable Contextual Multi-Armed Bandit (CMAB) problem over a cognitive state space $\mathcal{C}$, with theoretical grounding in Lipschitz smooth rewards and sublinear regret. This foundation enables a meta-controller that dynamically triggers System 2 based on visitation counts and confidence scores, balancing exploration and uncertainty in real time.
3.1 Theoretical Foundation
To provide a principled understanding of how exploration can be efficiently managed in dialog POMDPs, we develop a theoretical analysis that frames the problem as CMAB over a structured cognitive state space $\mathcal{C}$. This formulation rests on three pragmatic foundations: (1) compression of the high dimensional belief state into a low dimensional cognitive representation; (2) a Lipschitz smoothness assumption on the reward function; and (3) a derived bandit style exploration criterion. Together, these steps yield a tractable foundation that directly informs the design of our meta-controller.
3.1.1 CMAB Formulation
We formulate dialog policy learning as a CMAB problem (Foster and Rakhlin, 2020) to render the exploration-exploitation trade-off analytically tractable. The core innovation is a structured cognitive state space $\mathcal{C}$ that compresses the high dimensional belief state into a low dimensional and interpretable representation, thereby bridging bandit theory with dialog POMDPs.
In this CMAB, the arms are $\mathcal{A}={\text{S1},\text{S2}}$ (fast inference System 1 vs. deliberative reasoning System 2). The context is the cognitive state $\mathbf{c}_{t}=[d_{t},u_{t},\rho_{t}]\in\mathcal{C}$ (Computation details in Appendix A.1), which captures dialog progress, user uncertainty, and slot dependency. The reward $r_{t}(a)$ measures task progress and efficiency when choosing arm $a$ in context $\mathbf{c}t$. The objective is to minimize cumulative regret, where $a^{*}_{t}$ is the optimal arm and $a_{t}$ is the chosen arm at turn $t$.
$R_{T}=\sum_{t=1}^{T}\left[\mathbb{E}[r_{t}(a^{*}_{t}\mid\mathbf{c}_{t})]-\mathbb{E}[r_{t}(a_{t}\mid\mathbf{c}_{t})]\right].$ (1)
This formulation treats S2 as an oracle arm. When it invoked, aggressively pursues the optimal action in under explored regions. This directly informs the design of our meta-controller (Sec. 3.2.3).
3.1.2 Reward Smoothness Assumption
To support principled exploration in the cognitive state space $\mathcal{C}$, we assume the reward function is Lipschitz continuous (Asadi et al., 2018; Pazis and Parr, 2013; Ortner and Ryabko, 2012). This standard regularity condition ensures that nearby cognitive states yield similar rewards.
Assumption 3.1 (Lipschitz Smooth Reward in $\mathcal{C}$).
The expected immediate reward $\bar{r}(\mathbf{c},a)=\mathbb{E}[r(s_{t},a_{t})|\mathbf{c}_{t}=\mathbf{c}]$ is Lipschitz continuous with respect to the cognitive state $\mathbf{c}$ for any action $a$. That is, there exists a constant $L_{r}>0$ such that:
$|\bar{r}(\mathbf{c},a)-\bar{r}(\mathbf{c}^{\prime},a)|\leq L_{r}\cdot d(\mathbf{c},\mathbf{c}^{\prime}),\quad\forall\mathbf{c},\mathbf{c}^{\prime}\in\mathcal{C}.$
This assumption makes visitation counts meaningful proxies for epistemic uncertainty and allows bandit style exploration guarantees to carry over to the dialog POMDP. We empirically validate its practical relevance in Sec. 5.3.
Figure 2: The DyBBT Architecture. A meta-controller uses the cognitive state $\mathbf{c}_{t}$, visitation count $n_{t}(\mathbf{c}_{t})$, and System 1’s confidence $p_{t}^{S1}$ to dynamically select between System 1 (fast intuitive) and System 2 (slow deliberative). Outputs drive action execution and update visitation/distillation buffers for continuous learning.
3.1.3 Dynamic Balance Principle
Building on Assumption 3.1, we derive a principled exploration criterion for the meta-controller by formalizing the exploration-exploitation trade-off as a contextual bandit problem (Kleinberg et al., 2008; Bubeck et al., 2011) The exploration bonus for cognitive state $\mathbf{c}_{t}$ is:
$\text{Exploration-Bonus}(t)\propto\sqrt{\frac{\log T}{n_{t}(\mathbf{c}_{t})}},$ (2)
where $T$ is the total training steps and $n_{t}(\mathbf{c}_{t})$ is the visitation count. This adapts the Upper Confidence Bound (UCB) principle (Ortner and Ryabko, 2012; Foster and Rakhlin, 2020) to the structured cognitive space, with the square root and logarithmic terms arising from concentration inequalities and horizon scaling (Komiyama et al., 2024).
Eq. 2 motivates our meta-controller design. S2 is triggered when the exploration bonus exceeds a threshold, leading to Condition 1: $n_{t}(\mathbf{c}_{t})<\tau\sqrt{\log T}$. The scalar $\tau$ adjusts the trade-off, analogous to the confidence radius in UCB algorithms.
Under Assumption 3.1 and the approximate MDP structure in $\mathcal{C}$, this exploration strategy achieves sublinear regret (Proof sketch in Appendix A.2), demonstrating efficient and principled exploration in the compressed cognitive space.
3.2 System Architecture
Building on the theoretical foundation, DyBBT as shown in Fig. 2, operationalizes the CMAB formulation over the cognitive state space $\mathcal{C}$ into a dual-system architecture. The meta-controller directly instantiates the bandit-inspired switching rule (Eq. 2) to dynamically balance between fast intuitive S1 and slow deliberative S2. This principled design ensures expensive S2 is invoked only when cognitive signals and visitation counts indicate high epistemic uncertainty or low confidence, achieving adaptive exploration-exploitation trade-off while maintaining computational efficiency.
3.2.1 S1: Fast Intuitive Inference
S1 serves as the low latency, high throughput policy for most dialog turns, avoiding the prohibitive cost of perpetual deliberation. Given the current belief state, S1 outputs both a system action $a_{t}^{S1}$ formalized as a tuple $\{actiontype,domain,slot\}$, and a confidence score $p_{t}^{S1}\in[0,1]$. This score captures aleatoric uncertainty, complementing the epistemic uncertainty tracked by the meta-controller. S1 is trained in two stages: supervised fine-tuning on expert trajectories to predict actions and calibrated confidence, followed by PPO optimization for task success and efficiency.
3.2.2 S2: Slow Deliberative Reasoner
S2 is invoked only for novel or complex states where S1 is likely to fail. It uses the same frozen base model as S1 to retain broad knowledge. Upon activation, S2 generates three distinct action sequences, evaluates each based on the ratio of filled key slots, and selects the first action of the highest rated sequence as $a_{t}^{S2}$. Although computationally expensive, S2 provides robust reasoning in high uncertainty or high stakes scenarios, as guided by the meta-controller.
3.2.3 Meta-Controller
The meta-controller implements the bandit inspired exploration criterion (Eq. 2) by dynamically selecting between S1 and S2 based on real time cognitive signals. This dual-trigger mechanism bridges bandit theory with practical dialog POMDPs:
Activate S2 If:
$\underbrace{n_{t}(\mathbf{c}_{t})<\tau\sqrt{\log T}}_{\text{Condition 1}}\;\lor\;\underbrace{p_{t}^{S1}<\kappa}_{\text{Condition 2}}.$ (3)
Condition 1: Exploration. This condition directly implements the exploration bonus from Eq. 2. Under Assumption 3.1, low visitation counts in cognitive region $\mathbf{c}_{t}$ indicate high epistemic uncertainty, justifying systematic exploration via S2. The threshold $\tau\sqrt{\log T}$ adapts the classical bandit confidence radius to our structured cognitive space, ensuring exploration occurs when potential information gain outweighs computational cost.
Condition 2: Confidence. This condition addresses aleatoric uncertainty from partial observability and model limitations. Empirical studies show LLM confidence scores correlate with calibration (Kadavath et al., 2022; Lin et al., 2022; Yin et al., 2023); thus $p_{t}^{S1}<\kappa$ triggers S2 when S1’s parametric knowledge is likely insufficient.
This hybrid design ensures that S2 is invoked either for systematic exploration or as a robustness safeguard. The disjunctive combination yields an adaptive balance that outperforms either condition alone, as validated in our ablation study (Table 2). The meta-controller’s decisions form a closed loop system: high quality demonstrations from S2 are distilled into S1 via knowledge distillation (Appendix B.5.3), creating a virtuous cycle of policy improvement while progressively reducing long term reliance on costly deliberation.
4 Experiment
4.1 Experimental Setup
Datasets. We evaluate DyBBT on the Microsoft Dialog Challenge (Li et al., 2018) for single-domain tasks, and MultiWOZ 2.1 (Eric et al., 2020) for multi-domain tasks. Both are widely adopted in prior work. See Appendix B.1 for statistics.
Baselines. We compare DyBBT with a comprehensive set of recent and competitive baselines to ensure a rigorous evaluation, including previous SOTA EIERL. Full details are provided in Appendix B.2.
Evaluation Metrics. For single-domain tasks: success rate, average turns, and reward (following EIERL (Zhao et al., 2025): +2t+2t for success, t-t for failure, 1-1 for every turn). For multi-domain: Inform, Success, Book rates, and Avg. Turns (formulas in Appendix B.3).
Implementation Details. Following EIERL for fair comparison, dialogs are capped at 30 (single-domain) and 40 (multi-domain) turns. Training runs for 500 epochs (single) and 10K epochs (multi). DyBBT uses the same Qwen3 (0.6B-8B) for both S1 and S2. Full details in Appendix B.5.
4.2 Main Results Analysis
[…]
Table 1: Evaluation results for all agents across the three single-domain datasets are provided, with the highest value in each metric column highlighted in bold. Epochs (50, 250, 500) represent early, mid, and post convergence training stages. Baselines sourced from Zhao et al. (2025).
(a) Movie
(b) Rest
(c) Taxi
Figure 3: Learning curves for training efficiency and convergence across single-domain TODS tasks.
4.2.1 Performance on single-domain Tasks
The results in Table 1 show that DyBBT achieves strong performance across all three single-domain tasks. DyBBT’s cognitive state representation enables more efficient policy learning, leading to higher success rates with fewer turns compared to baselines. This advantage is especially pronounced in complex domains like Taxi, where slot dependencies create challenging exploration landscapes that DyBBT navigates effectively through its principled switching mechanism.
4.2.2 Performance on multi-domain Tasks
Appendix Table 6 presents results on the challenging MultiWOZ dataset. DyBBT maintains strong performance, whereas EIERL’s success rate drops significantly, revealing scalability limits of its population based approach. DyBBT-8B outperforms AutoTOD and ProTOD, and using GPT-4 as S2 achieves SOTA results, demonstrating that DyBBT matches strong LLM baselines with greater efficiency. This is enabled by the structured cognitive state and dual-system design, which provide domain-agnostic inductive bias without task-specific tuning. Cost-effectiveness analysis is discussed in Appendix E.7.
4.2.3 Training Efficiency and Convergence
Fig. 3 illustrates that DyBBT converges faster and achieves higher asymptotic performance than baselines across all domains, significantly outperforming EIERL as early as epoch 50. This accelerated learning stems from the meta-controller’s active guidance, which systematically targets under explored or uncertain regions in $\mathcal{C}$ instead of relying on random or high variance exploration. DyBBT also scales consistently with model size: success rates improve from 80.3% to 89.5% in the single-domain Movie task and from 78.2% to 84.1% in multi-domain settings when scaling from 0.6B to 8B. This indicates that the dual-system architecture effectively harnesses model ability. Coupled with efficient resource allocation via the meta-controller and Qwen3’s native switching mechanism (Appendix E.8), DyBBT demonstrates practical viability for real world deployment.
4.2.4 Key Advantages of DyBBT
The main results demonstrate that DyBBT achieves SOTA performance through the following key advantages: Dynamic Exploration Exploitation Balance: The bandit-inspired switching rule of the meta-controller enables DyBBT to allocate expensive S2 reasoning only when necessary, leading to highly efficient exploration. Scalability with Model Size: DyBBT benefits predictably from larger backbone models, making it well suited for future advances in LLM abilities. Strong Generalization: Consistent performance across both single- and multi-domain tasks shows that the cognitive state representation captures universal dialog dynamics. Computational Practicality: DyBBT maintains moderate computational overhead during both training and inference, unlike population based methods or full GPT-4.0 approaches.
4.3 Ablation Experiment
Table 2: Ablation study of DyBBT’s components on MultiWOZ. Results underscore the necessity of the meta-controller and the structured cognitive state representation for optimal performance.
[…]
Ablation results are shown in Table 2, and detailed settings are in Appendix E.2. The results reveal that: Meta-Controller is crucial. Removing it causes the most severe performance degradation, confirming its essential role in dynamically orchestrating the exploration-exploitation trade-off. Both conditions are necessary but asymmetric: Removing Condition 1 (EC) eliminates the bandit inspired exploration bonus from Eq. 2, while removing Condition 2 (CC) disables the aleatoric uncertainty safeguard, a distinction rooted in Bayesian RL theory (Dearden et al., 1998). Removing the confidence condition (CC) causes a more substantial performance drop than removing the exploration condition (EC), validating our hybrid design. This indicates that mitigating S1’s overconfidence is slightly more critical than targeted exploration for robust performance. In depth error analysis (Appendix E.3) reveals that CC primarily prevents catastrophic failures in states with high cognitive uncertainty. Cognitive State design is vital. Replacing it with the raw belief state causes catastrophic performance collapse, confirming the necessity of our low dimensional, interpretable representation. While the learned alternative performs reasonably well, it still underperforms our hand-designed features, justifying our cognitively inspired approach. All state dimensions contribute meaningfully. Removing any single dimension causes noticeable performance degradation, with dialog progress ($d_{t}$) being the most impactful individual component, followed by user uncertainty ($u_{t}$) and slot dependency ($\rho_{t}$). Knowledge Distillation enables continuous improvement. Disabling it reduces final performance, confirming its role in facilitating long term efficiency gains through systematic learning from S2’s demonstrations.
4.4 Human and Real World Evaluation
We conducted controlled human evaluations and real world user experiments to validate DyBBT’s practical efficacy beyond automated metrics.
Human Evaluation. Following the protocol in Appendix C, 10 NLP researchers evaluated 200 dialog states from MultiWOZ, comparing DyBBT against random switching and S1-only baselines. Annotators rated action appropriateness (5-point Likert scale) and judged whether invoking S2 was justified. DyBBT’s actions were rated as more appropriate than both baselines, and its decisions to invoke S2 aligned significantly better with human judgment than random switching, confirming that our meta-controller effectively identifies when deliberation is warranted, an affordance often missed by heuristic approaches.
Real World User Experiments. As detailed in Appendix D, 30 volunteers completed multi-domain dialog tasks. DyBBT achieved the highest task success rate and user satisfaction, demonstrating that its cognitive state representation $\mathcal{C}$ generalizes effectively beyond simulated environments. Case studies further showed that DyBBT successfully handles challenging scenarios, such as mid-dialog intent shifts and vague user expressions through adaptive S2 invocation.
Summary. These results provide converging evidence that DyBBT’s meta-controller translates cognitive affordances into a dynamic exploration exploitation balance, enabling robust performance in both controlled and real world settings.
5 Analysis
Our experimental results demonstrate that DyBBT achieves state-of-the-art performance on multiple benchmarks. In this section, we analyze the underlying mechanisms that enable DyBBT’s effectiveness, providing insights into why and how our framework works.
5.1 Analysis of Cognitive State Space
The cognitive state space $\mathcal{C}$ enables bandit style exploration in dialog POMDPs by providing a low dimensional and interpretable representation of dialog dynamics. To validate its structure, we analyze the visitation frequency across discretized regions of $\mathcal{C}$ during training (Fig. 5). The visitation heatmap reveals a structured, non-uniform pattern, confirming that exploration is guided by cognitive affordances: In early dialog phases ($d_{t}\in[0.0,0.2]$), the meta-controller broadly explores across high user uncertainty ($u_{t}$) for information gathering. In mid-phase ($d_{t}\in[0.4,0.6]$), visitation concentrates in medium-to-high $u_{t}$ regions to resolve ambiguities. In late phase ($d_{t}>0.8$), activity shifts to low $u_{t}$ states, focusing on exploitation to complete tasks.
This phase dependent targeting demonstrates that $\mathcal{C}$ effectively captures dialog progression and uncertainty, allowing the meta-controller to allocate exploration efficiently. The compactness and interpretability of $\mathcal{C}$ make principled exploration feasible in high dimensional dialog state spaces.
5.2 Adaptive Balancing and Continuous Improvement
The meta-controller’s hybrid triggering mechanism robustly addresses the exploration exploitation dilemma by responding to complementary forms of uncertainty. Condition 1 tackles epistemic uncertainty (lack of environmental knowledge), systematically exploring novel cognitive states, while Condition 2 addresses aleatoric uncertainty (inherent stochasticity or model limitations), acting as a consistent safety net against over reliance on a potentially flawed S1. Analysis of 10,000 dialog turns (Appendix Fig. 5) reveals their complementary temporal patterns: the exploration condition dominates early in training and for novel states, enabling systematic coverage, whereas the confidence condition provides ongoing robustness throughout training.
This dual trigger design naturally evolves with training progress. Initially, frequent S2 invocations provide guided exploration and high quality demonstrations. As S1 improves via knowledge distillation from these demonstrations, the meta-controller automatically reduces S2 usage, transitioning from guided exploration toward autonomous operation. This virtuous cycle enables continuous policy improvement without additional environment interactions. The distillation effectiveness is evidenced by monotonic improvement in S1 performance and corresponding reduction in S2 invocation rate. (Appendix Fig. 7). The adaptive balancing directly embodies the framework’s ability to perceive and respond to dynamic dialog affordances, ensuring appropriate cognitive resource allocation throughout the learning process.
5.3 Theoretical Intuitions and Empirical Alignment
Our theoretical analysis, though based on simplifying assumptions, is pragmatically validated by empirical results: Sublinear Regret as Validation of Core Assumptions. The empirical cumulative regret (Fig. 7) exhibits $\sqrt{T}$-like growth. This sublinear trend is not merely observational; it provides indirect empirical support for our key theoretical assumptions: The Lipschitz continuity of the reward in $\mathcal{C}$ (Assumption 3.1), and the approximate structure of MDP over $\mathcal{C}$ (Assumption A.1). The alignment between theory and experiment suggests $\mathcal{C}$ effectively captures the latent structure enabling efficient exploration. Low Dimensional $\mathcal{C}$ Enables Practical Implementation. The consistent high performance of DyBBT using only a three dimensional cognitive state demonstrates that the essential features governing exploration (dialog progress, user uncertainty, slot dependency) can be distilled into a compact representation. This reduction in dimensionality is theoretically motivated by the dependence of the regret bound’s $\sqrt{\dim(\mathcal{C})}$ (Appendix A.2.2).
5.4 Failure Mode Analysis
DyBBT exhibits three concrete failure modes empirically validated in Appendices E.9 and E.10. First, reliance on handcrafted cognitive state fidelity can lead to misrepresentation of complex dialog dynamics, causing the meta-controller to misjudge S2 invocation and resulting in underexploration or computational waste. Second, dependency on high quality S2 demonstrations introduces risk; errors in reasoning or self-evaluation can propagate to S1 via knowledge distillation, causing subtle policy corruption. Third, heuristic quantization of $\mathcal{C}$ into a fixed number of bins masks critical state variations, treating strategically distinct states identically and reducing exploration efficacy. Qualitative case studies illustrate how failures arise from unrepresented dialog nuances and how successful interventions align with human judgment, and reveals these failures affect only 5.2% of dialogs, primarily in edge cases with abrupt intent shifts or complex dependencies, while built-in safeguards provide substantial mitigation.
6 Conclusion
DyBBT presents a novel dialog policy learning framework that dynamically balances exploration and exploitation through a bandit inspired meta-controller grounded in a structured cognitive state space. By formalizing dialog affordances into interpretable dimensions: phasic progress, user uncertainty, and slot dependency, our method enables adaptive switching between fast intuitive responses and deliberate reasoning. Extensive experiments across single- and multi-domain benchmarks demonstrate SOTA performance in success rate, efficiency and generalization, with human evaluations confirming superior decision alignment. Future work will explore end-to-end learning of cognitive representations and extend the framework to more complex interactive settings.
Limitations
While DyBBT demonstrates strong empirical performance, its reliance on hand designed cognitive state representations, may not capture all nuances of highly complex or novel dialog dynamics. This points to a natural direction for future work: exploring end-to-end learning of cognitive representations that can adaptively refine the state space from interaction data, thereby extending the framework’s applicability to more diverse and intricate interactive settings.
Ethics Statement
This work presents a dialog policy learning framework evaluated on publicly available benchmark datasets (MS Dialog and MultiWOZ) and their use complies with the consent agreements established during their original release. These datasets have been previously anonymized and do not contain personally identifiable information or offensive content. Our research does not involve human subjects beyond the use of standard datasets, and all experiments are conducted through simulated user interactions. In the human evaluation and real user experiments, participants were volunteers proficient in English and knowledgeable in NLP, recruited from academic networks in Europe, the United States, and China. No monetary compensation was provided. Informed consent was obtained from all participants, and the study protocol followed established ethical guidelines for non-interventional research. The proposed methodology focuses on improving the efficiency of task-oriented dialog systems, with potential positive societal impacts through enhanced human computer interaction. We are unaware of any specific ethical concerns or negative social impacts directly arising from this work.
We utilized DeepSeek V3.2 for translation assistance and grammatical refinement of certain textual passages, and employed Qwen3-Code to aid in debugging and optimizing portions of the experimental code. These LLMs served solely as support tools for improving linguistic clarity and technical implementation. They played no role in the conceptualization of the research. The authors assume full responsibility for all aspects of the work, including the accuracy and integrity of all generated and modified content, and affirm that appropriate measures have been taken to prevent plagiarism and other forms of scientific misconduct.
Appendix A Theoretical Details
This section provides the theoretical motivation and intuition behind the DyBBT framework. The following analysis bridges ideas from bandit theory and cognitive science to create a heuristic for exploration in dialog POMDPs. While the full dialog POMDP problem is intractable for a rigorous minimax analysis, our goal is to provide a strong conceptual foundation and explanatory power for the algorithm’s design, which is then validated empirically in the main text.
A.1 Formalization of Cognitive State Space
[…]
The choice of these three dimensions is motivated by their central role in governing the exploration-exploitation trade-off in TODS, drawing inspiration from cognitive science and dialog theory:
  • Dialog Progress ($d_{t}=t/L$) captures the temporal affordance. Early phases ($d_{t}\rightarrow 0$) inherently afford more exploration to gather information, while late phases ($d_{t}\rightarrow 1$) afford exploitation to complete the task. This aligns with the common practice of annealing exploration schedules but provides a continuous, state dependent signal.
  • User Uncertainty operationalizes the information gathering affordance.
    $u_{t}=|S_{unconfirmed}|/|S_{relevant}|$
    A high $u_{t}$ indicates ambiguity in the user’s goal, directly signaling the need for information seeking actions to reduce entropy, a well established principle in decision theory.
  • Slot Dependency captures the structural affordance of the task environment, derived from a pre-computed slot co-occurrence matrix $M$ from the training corpus.
    $\rho_{t}=\max_{u\in U}(\frac{1}{|F|}\sum_{f\in F}M(u,f))$
    A high $\rho_{t}$ suggests that the next piece of information is highly predictable given what is already known (e.g., requesting departure after knowing destination in a taxi domain), making targeted exploitation more efficient than random exploration. This dimension encodes the latent structure of the domain.
This design transforms the complex, unstructured exploration problem in the raw belief space into a more manageable one in a structured space where states with similar exploration needs are grouped together, as visualized in Fig. 5.
A.2 Regret Analysis Under Simplifying Assumptions
To provide theoretical intuition for our exploration principle, we present a regret analysis under a set of simplifying assumptions that capture the core structure that we aim to exploit. This analysis justifies the form of our exploration bonus and provides an upper bound on learning speed. We make the following assumptions to bridge the gap between bandit theory and the dialog POMDP. Our analysis is based on the Assumption 3.1 stated in Sec. 3.1.2, which posits Lipschitz smoothness of the reward function in the cognitive state space $\mathcal{C}$.
Assumption A.1 (MDP over $\mathcal{C}$).
The dialog process can be approximately modeled as a finite horizon MDP over the cognitive state space $\mathcal{C}$. The transition dynamics and expected reward $\bar{r}(\mathbf{c},a)=\mathbb{E}[r(s_{t},a_{t})|\mathbf{c}_{t}=\mathbf{c}]$ depend primarily on $\mathbf{c}_{t}$.
The value function under a policy $\pi$ in the cognitive state space is defined as:
$V^{\pi}(\mathbf{c})=\mathbb{E}\!\left[\sum_{k=0}^{H}\gamma^{k}\,\bar{r}(\mathbf{c}_{t+k},a_{t+k})\;\middle|\;\mathbf{c}_{t}=\mathbf{c},\right.\\ \left.a_{t+k}\sim\pi(\cdot|\mathbf{c}_{t+k})\right].$ (4)
This assumption is a pragmatic simplification that allows us to focus on the core exploration challenge. It is reasonable if the cognitive state $\mathbf{c}_{t}$ is a sufficient statistic for the exploration-exploitation trade-off, which our empirical results support.
A.2.1 Theoretical Intuition for Regret
Under Assumptions 3.1 and A.1, if we perform optimistic exploration in the cognitive state space $\mathcal{C}$, prioritizing states with low visitation counts, we can derive an upper bound on the expected cumulative regret that scales sublinearly with time:
$\mathbb{E}[R(T)]\lesssim\widetilde{\mathcal{O}}\left(L_{r}\cdot\sqrt{\dim(\mathcal{C})\cdot T}\right),$ (5)
where $R(T)=\sum_{t=1}^{T}[V^{*}(\mathbf{c}_{t})-V^{\pi_{t}}(\mathbf{c}_{t})]$ is the cumulative regret, and $\widetilde{\mathcal{O}}$ hides logarithmic factors. The notation $\lesssim$ indicates that this is a heuristic bound that captures the expected asymptotic scaling rather than a rigorous inequality. Here, $L_{r}$ is the Lipschitz constant from Assumption 3.1, bounding the reward’s sensitivity to changes in $\mathcal{C}$.
A.2.2 Derivation Sketch
This scaling can be motivated by discretizing the cognitive state space $\mathcal{C}$ into $N=\mathcal{O}((1/\epsilon)^{\dim(\mathcal{C})})$ cells of diameter $\epsilon$.
  1. 1.
    Discretization Error: Due to Lipschitz continuity of $\bar{r}(\mathbf{c},a)$ (Assumption 3.1), the error introduced by discretization is bounded by $\mathcal{O}(L_{r}\epsilon T)$.
  2. 2.
    Bandit Regret: For the discretized MDP with $N$ state cells, treating each cell arm analogously, a UCB like algorithm can achieve a regret bound of $\mathcal{O}(\sqrt{NT\log T})$.
  3. 3.
    Optimization: Balancing the two error terms by setting $\epsilon\sim T^{-1/(\dim(\mathcal{C})+2)}$ yields the final bound $\widetilde{\mathcal{O}}(L_{r}\cdot\sqrt{\dim(\mathcal{C})\cdot T})$.
This sketch illustrates that efficient learning is possible by exploiting the low dimensional structure and smoothness of the value function in $\mathcal{C}$, providing intuition for our exploration criterion.
This bound provides an intuitive justification for our exploration criterion (Eq. 2 in the main text). The term $\sqrt{\frac{\log T}{n_{t}(\mathbf{c}_{t})}}$ is a heuristic adaptation of the optimism principle, encouraging exploration of states with high uncertainty, inversely proportional to their visitation count. The empirical regret curve (Fig. 7) shows sublinear growth, consistent with this theoretical intuition.
A.3 Justification for the Meta-Controller Rule
The meta-controller’s hybrid rule is designed for robust performance in the realistic setting where our theoretical assumptions hold only approximately:
Activate S2 IF:
$\left(n_{t}(\mathbf{c}_{t})<\tau\sqrt{\log T}\right)\lor\left(p_{t}^{S1}<\kappa\right).$ (6)
The first condition, $n_{t}(\mathbf{c}_{t})<\tau\sqrt{\log T}$, is the direct implementation of the theoretical exploration principle derived above. It addresses epistemic uncertainty ( uncertainty reducible by exploration) by triggering System 2 in regions of $\mathcal{C}$ that are under explored relative to the time horizon.
The second condition, $p_{t}^{S1}<\kappa$, is a critical empirical safeguard that addresses limitations of the theoretical model:
  • Partial Observability: The true state of the user may not be fully captured by the belief state $\mathbf{s}_{t}$, leading to aleatoric uncertainty.
  • Model Imperfection: S1, as a parameterized policy, may have inherent limitations and blind spots not captured by the visitation count.
  • Assumption Violation: The Lipschitz smoothness assumption may locally break down.
A low confidence score $p_{t}^{S1}$ is a proxy for these forms of uncertainty. This condition ensures robustness by invoking the powerful, knowledge rich S2 when S1 is uncertain, preventing catastrophic failures. The disjunctive ($\lor$) combination ensures System 2 is activated for either theoretical exploration or empirical robustness, making the overall system more adaptive and reliable than either condition alone, as evidenced by the ablation study (Table 2).
A.4 Discussion and Limitations
Our theoretical analysis provides a formal motivation for the DyBBT framework by illustrating how exploiting the structure of a cognitive state space can lead to efficient exploration. However, we acknowledge its limitations, which also highlight the value of our empirical validation:
Simplified Model: Assumption A.1 reduces the POMDP to an MDP over $\mathcal{C}$, ignoring the challenges of belief state tracking and partial observability. This is a significant simplification. Our empirical results show that the algorithm performs well even when this assumption is not perfectly met, as the meta-controller’s confidence condition can mitigate some of these issues.
Heuristic Adaptation: The exploration bonus and the meta-controller rule are heuristic adaptations of the theoretical principle. A rigorous derivation for POMDPs remains an open challenge. Our contribution is to demonstrate that this heuristic is well motivated and highly effective in practice.
Empirical Safeguard: The confidence based condition, while crucial for performance, is not derived from the regret analysis. Its justification is empirical, stemming from its necessity for robust performance in ablation studies.
In conclusion, the theoretical analysis is not intended as a strict performance guarantee but rather as an explanatory framework that provides strong intuition for why exploring based on cognitive state visitation counts is a powerful principle. The ultimate validation of this principle, and its pragmatic implementation in the meta-controller, lies in its consistent empirical success across diverse dialog benchmarks.
Appendix B Experiment Details
B.1 Experimental Platform and Datasets
We evaluated DyBBT on two widely adopted benchmarks: the Microsoft dialog Challenge (MS dialog) (Li et al. (2018)) for single-domain tasks, and the MultiWOZ 2.1 corpus (Eric et al. (2020)) for multi-domain tasks. Both datasets are converted into ConvLab-3’s unified format, ensuring consistency in ontology, state representation, and API interaction. Table 3 summarizes the key statistics of both datasets.
The MS Dialog dataset comprises three distinct domains: Movie-Ticket Booking, Restaurant Reservation, and Taxi Ordering. It contains 7,215 dialogs with 89,465 turns, averaging 12.4 turns per dialog. The dataset is partitioned into training, validation, and test sets with 5,772, 722, and 721 dialogs, respectively.
The MultiWOZ 2.1 dataset is a large scale multi-domain corpus spanning seven domains: Attraction, Hotel, Restaurant, Taxi, Train, Hospital, and Police. It includes 10,420 dialogs and 145,360 turns, with an average of 13.9 turns per dialog. The dataset is split into 8,420 dialogs for training, 1,000 for validation, and 1,000 for testing.
Both datasets provide annotated belief states, system dialog acts, and user goals, making them suitable for training and evaluating end-to-end dialog policies. The diversity in domain complexity, dialog length, and task structure across these datasets allows us to thoroughly assess the generalization capability of DyBBT in both single and multi-domain settings.
To ensure reproducibility and enable fair comparison, we implement and evaluate our proposed DyBBT framework using ConvLab-3 (Zhu et al., 2023), a flexible and unified toolkit for TODS. ConvLab-3 provides standardized data formats, integrated user simulators, and reinforcement learning utilities, facilitating consistent development and evaluation of dialog policies across multiple domains. All experiments are conducted using ConvLab-3’s builtin simulators and evaluation metrics, ensuring comparability across models and domains.
Dataset Domains Dialogs Turns Avg.Turns/Dialog
MS Dialog 3 7,215 89,465 12.4
MultiWOZ 2.1 7 10,420 145,360 14.0
Table 3: Summary of dataset statistics for MS Dialog and MultiWOZ 2.1.
B.2 Baselines Details
[…]
  • NOISY_\_DQN agents enhance exploration by introducing noise into the network weights, based on the stable noisy network (NROWAN-DQN) with noise reduction and online weight adjustment (Han et al., 2022).
  • PG (REINFORCE) is a stochastic gradient algorithm for policy gradient reinforcement learning, and its implementation refers to the flexible dialog system toolkit ConvLab-3 to serve as a dialog policy baseline (Zhu et al., 2023).
  • PPO is a policy optimization method in policy-based reinforcement learning that uses multiple epochs of stochastic gradient ascent and a constant clipping mechanism as the soft constraint for each policy update, with its implementation relying on the ConvLab-3 dialog toolkit (Zhu et al., 2023).
  • LLM_\_DP agents replace the dialog policy (DP) module of the TODS with GPT-4.0 (drawing on advances in LLM based multi turn dialog systems) to select appropriate actions and pass them to the natural language generation (NLG) module for response generation (Yi et al., 2025).
  • AutoTOD is a zero-shot autonomous agent based on GPT-4.0, which rethinks TODS by shifting from complex modularity to zero-shot autonomy and acts as a dialog policy baseline (Xu et al., 2024).
  • ProTOD is a proactive TODS policy based on GPT-4.0, designed as a proactive dialog system to optimize the process of task oriented interactions (Dong et al., 2025).
  • EIERL is an evolutionary reinforcement learning method for TODS policies, which improves the efficiency of dialog policy learning by injecting elite individuals into the evolutionary process (Zhao et al., 2025).
  • MACRM is a multi agent curiosity reward model for TODS, which optimizes dialog policies through collaborative interactions among multiple agents and curiosity driven reward mechanisms (Sun et al., 2025).
B.3 Metrics Formula
This section provides the formal definitions of the evaluation metrics used for multi-domain TODS evaluation, following the standard MultiWOZ evaluation protocol.
B.3.1 Inform Success Rate
The Inform Success Rate measures the system’s ability to provide all requested information to the user. Let $G$ be the goal specification, $D$ be the set of dialog domains, and $S$ be the sequence of system dialog acts. For each domain $d\in D$, let $R_{d}$ be the set of requested slots in the goal:
$\text{TP}=\sum_{d\in D}\sum_{s\in R_{d}}\mathbb{I}\left(\exists\,\text{inform}(d,s,v)\right.\\ \left.\in S\land v\notin V_{\text{null}}\right)$ (7)
$\text{FP}=\sum_{d\in D}\sum_{s\notin R_{d}\cup I_{d}}\mathbb{I}\left(\exists\,\text{inform}(d,s,v)\right.\\ \left.\in S\land v\notin V_{\text{null}}\right)$ (8)
$\text{FN}=\sum_{d\in D}\sum_{s\in R_{d}}\mathbb{I}\left(\nexists\,\text{inform}(d,s,v)\right.\\ \left.\in S\lor v\in V_{\text{null}}\right)$ (9)
where $V_{\text{null}}=\{``^{\prime\prime},\text{``dont care''},\text{``not mentioned''}\}$ represents null values. The Inform Success Rate is then defined as:
$\text{Inform}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (10)
B.3.2 Book Success Rate
The Book Success Rate evaluates the system’s ability to successfully complete booking operations. For each domain $d\in D$ that requires booking, let $B_{d}$ be the set of booking constraints in the goal. The booking success is computed as:
$\text{Book}_{d}=\frac{1}{|B_{d}|}\sum_{b\in B_{d}}\mathbb{I}\left(\text{book}(d,b,v)\right.\\ \left.\in S\land v=v_{\text{goal}}\right)$ (11)
For the taxi domain (which has no database constraints), booking success is trivially 1 if any booking action occurs:
$\text{Book}_{\text{taxi}}=\mathbb{I}\left(\exists\,\text{book}(\text{taxi},\cdot,\cdot)\in S\right)$ (12)
The overall Book Success Rate is the average across all booking domains:
$\text{Book}=\frac{1}{|D_{\text{book}}|}\sum_{d\in D_{\text{book}}}\text{Book}_{d}$ (13)
where $D_{\text{book}}$ is the set of domains requiring booking.
B.3.3 Success Rate
The Success Rate represents the overall task completion performance, combining both information provision and booking success:
$\text{Success}=\mathbb{I}\left(\text{Inform}=1\land\text{Book}=1\right)$ (14)
This binary metric indicates whether both all requested information was provided and all booking operations were successfully completed.
This metric rewards systems that achieve high success rates with fewer dialog turns, promoting both effectiveness and efficiency.
B.4 Prompt for DyBBT and LLM-DP
This appendix provides the detailed prompts used for System 1 (intuitive controller) and System 2 (reasoning controller) in the DyBBT framework. The LLM_DP prompt is the same from the EIERL paper(Zhao et al. (2025)).
B.4.1 System 1 Prompt
B.4.2 System 2 Prompt
B.4.3 LLM_DP Prompt
B.5 Implementation Details
The DyBBT framework was implemented within the Convlab-3 dialog system with Python 3.10 environment (Zhu et al. (2023)), leveraging its modular architecture for efficient dialog policy optimization. We employed RuleDST for system dialog state tracking and RulePolicy for user policy simulation, eliminating the need for natural language understanding (NLU) and natural language generation (NLG) modules. This design choice significantly enhances training efficiency by reducing computational overhead and isolating the impact of language processing components from policy learning performance. The dialog environment was configured with a maximum turn limit of 30 for single-domain and 40 for multi-domain (the same as EIERL) interactions per episode, with the cognitive state space $\mathcal{C}$ computed in real-time during dialog execution using dimensions including dialog progress ($d_{t}$), user uncertainty ($u_{t}$), and slot dependency ($\rho_{t}$) extracted from the belief state representation provided by RuleDST.
User goals were dynamically generated using the GoalGenerator module, which produces diverse and realistic TODS objectives across single or multiple domains. This approach ensures training data variety and generalization capability, consistent with REINFORCE and PPO training methodologies. The goal generation process excluded the police domain due to its low data quality, ensuring higher reliability in evaluation.
All experiments were conducted on NVIDIA 5090 GPUs with 32GB memory. System 1 was SFT using the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and further optimized via PPO, employing a clipping parameter $\epsilon=0.2$ and GAE with $\lambda=0.95$. The meta-controller employs a dual-threshold mechanism for System 2 invocation, with $kappa=0.7$ and $\tau=1.0$, values selected via grid search over development sets as they maximize both performance and robustness across domains. These thresholds operate on a discretized 5 bins cognitive state space, which balances expressiveness and generalization, as validated in Sec. E.5.
We maintained a replay buffer with a capacity of 10,000 transitions, using a batch size of 32 for training. A separate knowledge distillation buffer was managed under a FIFO replacement policy with a fixed capacity. To ensure reproducibility, all experiments were run with five fixed random seeds (9841, 35741, 91324, 8134, 13924), consistent with the EIERL baseline (Zhao et al., 2025). All hyperparameters were selected through grid search on a validation subset of the MultiWOZ data.
Training was conducted for 500 epochs on single-domain tasks and 10,000 epochs on multi-domain tasks, incorporating early stopping with a patience of 3 epochs based on validation performance. This protocol aligns with the EIERL setup for fair comparison.
B.5.1 Slot Co-occurrence Matrix Construction
The slot dependency dimension $\rho_{t}$ in the cognitive state space $\mathcal{C}$ is derived from a co-occurrence matrix $M$ that captures statistical relationships between dialog slots across the Microsoft dialog Challenge (Li et al. (2018)) and MultiWOZ (Eric et al. (2020)) dataset. This matrix quantifies the conditional probability that slot $j$ appears given the presence of slot $i$, providing a principled measure of semantic relatedness between dialog concepts.
Formally, the co-occurrence matrix $M\in\mathbb{R}^{N\times N}$ is constructed from the training partition of MultiWOZ 2.1, where $N$ represents the total number of unique slot types across all domains. For each dialog turn containing belief state updates, we extract the set of active slots (those with non-empty values) and update the co-occurrence counts. The matrix elements are computed as:
$M_{ij}=\frac{\text{count}(\text{slot}_{i}\land\text{slot}_{j})}{\text{count}(\text{slot}_{i})}$ (15)
where $\text{count}(\text{slot}_{i}\land\text{slot}_{j})$ denotes the number of dialog turns where both slots appear simultaneously, and $\text{count}(\text{slot}_{i})$ represents the total occurrences of slot $i$. This normalization ensures that $M_{ij}$ represents the empirical conditional probability $P(\text{slot}_{j}|\text{slot}_{i})$.
The slot dependency $\rho_{t}$ for a given belief state $s_{t}$ is then computed as the average co-occurrence strength between the currently active slots:
$\rho_{t}=\frac{1}{|A_{t}|(|A_{t}|-1)}\sum_{i\in A_{t}}\sum_{j\in A_{t},j\neq i}M_{ij}$ (16)
where $A_{t}$ denotes the set of slots with non-empty values in the current belief state. This formulation captures the structural complexity of the dialog context, with higher values indicating greater semantic interdependence between the information being discussed.
The construction of $M$ leverages the statistical regularities present in TODS, where certain slot combinations naturally co-occur due to domain-specific constraints and user behavior patterns. For instance, in restaurant booking scenarios, slots like restaurant-area and restaurant-food frequently appear together, while in hotel domains, hotel-pricerange and hotel-type exhibit strong associations. This matrix based approach provides a data-driven foundation for quantifying dialog complexity that complements the theoretically motivated dimensions of dialog progress and user uncertainty.
B.5.2 Training Details For System 1
To train System 1 for accurate action prediction and confidence estimation, we employ a two-stage training methodology comprising supervised fine-tuning (SFT) followed by reinforcement learning. This approach utilizes dialog sequences from the MultiWOZ and MS Diag dataset to develop a robust policy model capable of rapid decision making with calibrated confidence scores.
Stage 1: Supervised Fine-tuning with Data Augmentation
We first construct a training corpus of 10,000 single turn dialog samples through systematic data augmentation. For each dialog turn, we extract the belief state $\mathbf{s}_{t}$, available action set $\mathcal{SA}$, and ground truth system actions $a_{t}^{*}$. The initial confidence score $p_{t}^{S1}$ is sampled from $\mathcal{U}(0.95,1.0)$.
The augmentation process introduces controlled perturbations to simulate prediction uncertainty. For each ground truth action sequence $a_{t}^{*}$, we apply three modification operations with specified probabilities: 20% action addition by sampling new actions from $\mathcal{SA}$; 60% action modification through substitution with random actions from $\mathcal{SA}$; and 20% action deletion while ensuring the augmented sequence $a_{t}^{\prime}$ maintains at least one action. These operations are applied sequentially in random order to each sample (Kadavath et al., 2022; Lin et al., 2022; Yin et al., 2023). The confidence score is adjusted proportionally to the modification intensity:
$p_{t}^{S1}\leftarrow p_{t}^{S1}\cdot\left(1-\frac{n_{\text{mod}}}{n}\right),$
where $n$ denotes the original action sequence length and $n_{\text{mod}}$ represents the number of modified actions. This procedure generates a dataset with confidence scores approximately uniformly distributed in $[0,1]$.
For SFT training, the model takes $\mathbf{s}_{t}$ and $\mathcal{SA}$ as inputs and produces both action sequence $a_{t}^{S1}$ and confidence score $p_{t}^{S1}$ as outputs. The composite loss function integrates action prediction and confidence estimation:
$\mathcal{L}=\lambda\mathcal{L}_{\text{a}}+(1-\lambda)\mathcal{L}_{\text{p}},$
where $\lambda=0.7$. The action loss $\mathcal{L}_{\text{a}}$ employs cross-entropy to measure divergence between predicted and augmented actions:
$\mathcal{L}_{\text{a}}=-\sum_{i}\log P(a_{t}^{S1}=a_{t}^{\prime}\mid\mathbf{s}_{t},\mathcal{SA}),$
while the confidence loss $\mathcal{L}_{\text{p}}$ utilizes mean squared error:
$\mathcal{L}_{\text{p}}=\left(p_{t}^{S1}-p_{t}^{\text{target}}\right)^{2}.$
Stage 2: Reinforcement Learning with PPO
The second stage employs PPO to optimize dialog level performance metrics using the complete MultiWOZ dataset. The reward function $R$ combines multiple objectives:
$R=R_{\text{success}}+R_{\text{efficiency}}+R_{\text{penalty}},$
where $R_{\text{success}}=+2t$ for successful dialogs and t-t for failures ($t$ denotes the max turn number), $R_{\text{efficiency}}=-1$ per dialog turn to encourage conciseness, and $R_{\text{penalty}}$ captures additional constraints.
This two-stage approach enables System 1 to initially learn accurate action confidence mappings through supervised learning, then refine its policy for improved task completion efficiency and success rates via reinforcement learning.
B.5.3 Knowledge Distillation Buffer Management
To form a virtuous cycle and reduce long term dependence on System 2, high quality decisions $(s_{t},a_{t}^{S2})$ from System 2 are stored in a distillation buffer $D_{\text{distill}}$. We only store decisions where System 2’s self evaluated task completion probability is greater than 0.9, ensuring high quality distillation data. Periodically (every 10 training epochs), System 1 is fine-tuned on these data via Low-Rank Adaptation (LoRA) with a learning rate of $1\times 10^{-4}$, batch size of 4, and gradient accumulation steps of 8. This SFT approach distills the knowledge gained through costly deliberation into an efficient intuitive policy while maintaining computational efficiency, leading to a monotonic performance improvement. Over time, this reduces the need to invoke System 2 for previously challenging states, thereby increasing overall efficiency.
The knowledge distillation buffer $D_{\text{distill}}$ stores high quality pairs $(s_{t},a_{t}^{S2})$ generated by System 2. The buffer has a maximum capacity and uses an FIFO policy to maintain data freshness and diversity. We employ LoRA fine-tuning with rank $r=16$, scaling parameter $\alpha=32$, and dropout rate of 0.1, targeting the query and value projection layers of the transformer architecture. This configuration achieves parameter efficiency while preserving the base model’s generalization capabilities.
Algorithm 1 Knowledge Distillation Buffer Update and Sampling
1:Buffer Update:
2:Input: Current belief state $s_{t}$, System 2 action $a_{t}^{S2}$, System 2 self evaluated confidence $p_{\text{self}}$
3:if $p_{\text{self}}>0.9$ then $\triangleright$ Only store high confidence actions
4:  if $|D_{\text{distill}}|<\text{MAX\_SIZE}$ then
5:   $D_{\text{distill}}.\text{append}((s_{t},a_{t}^{S2}))$
6:  else
7:   $D_{\text{distill}}.\text{pop\_front}()$ $\triangleright$ Remove oldest entry (FIFO)
8:   $D_{\text{distill}}.\text{append}((s_{t},a_{t}^{S2}))$
9:  end if
10:end if
11:System 1 Fine-tuning:
12:Input: System 1 model with LoRA adapters, buffer $D_{\text{distill}}$
13:Every 10 training epochs:
14:for $epoch=1$ to 11 do $\triangleright$ Fine-tune for 1 epoch
15:  for each batch sampled from $D_{\text{distill}}$ do
16:   Compute loss $\mathcal{L}=\text{CrossEntropy}(\text{System1}(s_{i}),a_{i})$
17:   Update LoRA adapter parameters via gradient descent
18:  end for
19:end for
B.5.4 Visitation Count of the Cognitive State Space
To compute the visitation count $n_{t}(\mathbf{c}_{t})$ for the continuous cognitive state space $\mathcal{C}$, we discretize each dimension of $\mathbf{c}_{t}=[d_{t},u_{t},\rho_{t}]$ into 5 uniformly spaced bins over the range $[0,1]$. The cognitive state is then mapped to a discrete tuple $(d_{\text{bin}},u_{\text{bin}},\rho_{\text{bin}})$, and $n_{t}(\mathbf{c}_{t})$ is the cumulative visitation count of that bin tuple.
This choice of dimensions is motivated by cognitive and dialog theory, which highlights stage, uncertainty, and structural relationships as key factors influencing decision making. By quantifying these environmental affordances into a structured cognitive state space $\mathcal{C}$, we create a formal bridge between Gibson’s ecological perception theory and practical dialog policy optimization. While not exhaustive, this representation aims to capture the most salient features for guiding exploration. Its empirical necessity and sufficiency are validated through ablation studies in Sec. 4.3. We define $\mathcal{C}$ as the cognitive state space, assumed to be a compact subset of $\mathbb{R}^{3}$ equipped with the Euclidean metric $d(\mathbf{c},\mathbf{c}^{\prime})$.
B.5.5 Calculation of Empirical Cumulative Regret
To empirically validate the theoretical intuition of sublinear regret growth under our simplifying assumptions, we compute the empirical cumulative regret $R_{\text{emp}}(T)$ during training, as shown in Fig. 7. The regret is defined as:
$R_{\text{emp}}(T)=\sum_{t=1}^{T}\left(V^{\pi^{*}}(\mathbf{s}_{t})-V^{\pi_{t}}(\mathbf{s}_{t})\right)$
where:
  • $T$ is the total number of dialog turns (training steps) up to the current point.
  • $\mathbf{s}_{t}$ is the belief state at turn $t$.
  • $V^{\pi_{t}}(\mathbf{s}_{t})$ is the actual discounted return obtained from state $\mathbf{s}_{t}$ under the current policy $\pi_{t}$ at training step $t$.
  • $V^{\pi^{*}}(\mathbf{s}_{t})$ is the value of the near-optimal policy $\pi^{*}$ at state $\mathbf{s}_{t}$.
Since the true optimal policy $\pi^{*}$ is unknown, we approximate it using a strong baseline policy the fully trained DyBBT-8B/GPT-4.0 model, which achieves SOTA performance on MultiWOZ. We assume this policy is sufficiently close to optimal for regret estimation purposes. For each state $\mathbf{s}_{t}$, we estimate $V^{\pi^{*}}(\mathbf{s}_{t})$ by running $\pi^{*}$ from $\mathbf{s}_{t}$ for multiple episodes and averaging the discounted returns. Actual episodic return is used from the current dialog episode as a proxy for $V^{\pi_{t}}(\mathbf{s}_{t})$. Although this is a coarse approximation, it is standard in episodic RL settings and sufficient to capture the regret trend.
$R_{\text{emp}}(T)$ is plotted against $T$ on a log-log scale to clearly visualize the sublinear growth trend. The theoretical upper bound $\widetilde{\mathcal{O}}(\sqrt{T})$ is plotted alongside for comparison. The constant factor in the theoretical bound is fit to the empirical curve in the early training phase to align the curves for illustrative purposes.
Appendix C Human Evaluation Details
This appendix provides comprehensive details of the human evaluation study described in Sec. 4.4. The study was designed to qualitatively assess the core contribution of the DyBBT framework: the intelligent, adaptive decision making of its meta-controller, beyond what is captured by automated metrics.
C.1 Annotation Protocol and Interface
Evaluators were presented with a structured web interface for each evaluation instance. Each instance consisted of a single dialog state (not a full dialog), sampled from the MultiWOZ test set. For a given state, the interface displayed the following information:
  • Dialog Context: The last user utterance and the last system action to provide conversational context.
  • Current Belief State ($\mathbf{s}_{t}$): A structured table showing all relevant slots for the domain(s), their values, and their confirmation status (e.g., confirmed, requested, None).
  • Cognitive State ($\mathbf{c}_{t}$): The numerical values for dialog progress ($d_{t}$), user uncertainty ($u_{t}$), and slot dependency ($\rho_{t}$).
  • System Action: The action chosen by the model for this state, presented in a structured format (e.g., [request, restaurant, area, “ "]).
  • System Variant: The name of the model variant that produced the action (DyBBT, S1-only, Random Switching). Variants were anonymized as ‘System A’, ‘System B’ during evaluation to avoid bias.
Evaluators were then asked to answer two questions based solely on the provided information:
  1. 1.
    Action Appropriateness: “How appropriate is the system’s chosen action given the current dialog state?” Rated on a 5 points Likert scale:
    1. 1.
      Very Inappropriate
    2. 2.
      Somewhat Inappropriate
    3. 3.
      Neutral
    4. 4.
      Somewhat Appropriate
    5. 5.
      Very Appropriate
  2. 2.
    Switching Judgment: “In this specific situation, would it be justified to invoke a powerful, but computationally expensive, reasoning module to choose the action?” Answered with Yes or No. This question was only shown for states where the evaluated model did not invoke System 2, to directly test if the meta-controller’s decision not to invoke aligned with human judgment.
C.2 Annotator Background and Training
We recruited 10 annotators, all of whom were graduate students or researchers with a background in natural language processing and familiarity with TODS. Prior to the evaluation, a mandatory 30 minutes training session was conducted. The session:
  • Explained the goal of the evaluation and the definition of key concepts (belief state, system actions, computational cost).
  • Walked through 5 example states that were not part of the evaluation set, discussing potential appropriate actions and reasoning for/against invoking a costly reasoner.
  • Allowed annotators to ask questions to resolve any ambiguities.
Annotators were compensated at a competitive hourly rate for their work.
C.3 Human Evaluation Results
Table 4: Complete Human Evaluation Results. The Action Appropriateness score is the average Likert score (1-5). The Switching Agreement is the percentage of states where the model’s decision to not invoke System 2 aligned with the majority of human annotators.
Model Variant Action Appropriateness $\uparrow$ Switching Agreement $\uparrow$ DyBBT-8B 4.31 $\pm$ 0.12 88.7% w/o Meta-Controller (Random) 3.72 $\pm$ 0.19 52.3% w/ S1-only 3.95 $\pm$ 0.15 w/o Exploration Condition (EC) 4.08 $\pm$ 0.14 75.4% w/o Confidence Condition (CC) 3.89 $\pm$ 0.16 81.2%
The results in Table 4 provide a detailed breakdown supporting the main findings:
  • Superior Decision Quality: The full DyBBT model yields a higher action appropriateness score than the ablated variants.
  • Value of the Meta-Controller: The random switching variant has the lowest scores, confirming that a naive switching strategy severely degrades decision quality and is not aligned with human judgment.
  • Complementary Role of Both Conditions: Removing either the Exploration Condition (EC) or the Confidence Condition (CC) leads to a drop in both appropriateness and agreement, with the CC being slightly more critical for action quality (preventing poor actions) and the EC being crucial for efficient switching (preventing unnecessary calls). This validates their hybrid design in the meta-controller.
C.4 Qualitative Analysis of Meta-Controller Decisions
To qualitatively validate the efficacy of the meta-controller’s switching mechanism beyond aggregate metrics, we present two contrasting case studies sampled from the MultiWOZ test set. These examples illustrate how DyBBT’s principled switching aligns with human judgment, in contrast to a naive baseline.
Case 1: High Agreement Example (DyBBT). The meta-controller correctly identified a state warranting costly deliberation due to high aleatoric uncertainty despite the cognitive state being well explored. The belief state, cognitive signals, and subsequent action were as follows.
Cognitive State: $d_{t}=0.3$ (early-stage), $u_{t}=0.8$ (high uncertainty), $\rho_{t}=0.6$. Meta-Controller Decision: System 1’s confidence was low ($p_{t}^{S1}=0.6<\kappa$), triggering System 2 via the confidence condition. System 2 performed a multi path reasoning and produced a confirm_all action sequence to disambiguate the user’s intent: confirm(restaurant, area) and confirm(restaurant, name). Annotators overwhelmingly rated this intervention as appropriate (Avg: 4.8/5) and agreed (90%) that invoking System 2 was justified. This case demonstrates the critical role of the confidence condition as a robustness safeguard against System 1’s inherent limitations in partially observable contexts. Case 2: Low Agreement Example (Random Switching). A random switching baseline ($10\%$ chance per turn) invoked System 2 in a state where the optimal action was obvious, leading to computational waste without performance gain: Cognitive State: $d_{t}=0.9$ (late stage), $u_{t}=0.1$ (low uncertainty), $\rho_{t}=0.2$. Scenario: All user constraints are satisfied, and the booking is complete. The only appropriate action is to terminate the dialog with goodbye. The random controller invoked System 2, which also output goodbye. Annotators rated the action itself as appropriate (Avg: 4.2/5) but unanimously (100%) judged the invocation of System 2 as not justified, deeming it an inefficient use of resources. This highlights a key failure mode of static or non-adaptive switching heuristics and underscores the necessity of our cognitive state aware meta-controller. In summary, these cases provide concrete evidence that DyBBT’s switching mechanism dynamically allocates computational resources in a manner that is both effective and efficient, closely mirroring human expert judgment.
Appendix D Real World User Experiments
Table 5: Real World User Experiment Results. Success Rate measures the percentage of successfully completed dialogs. Average Turns counts the number of dialog turns per task. User Satisfaction is rated on a 1-5 Likert scale.
[…]
While all previous experiments relied on simulated users, real world user interactions are inherently more complex and unpredictable. This raises a key concern regarding generalization: user behavior in practice may not neatly align with the quantifiable dimensions of our cognitive state space $\mathcal{C}$, potentially limiting DyBBT’s applicability. To investigate this and verify the robustness of our assumptions, we conducted experiments with real human users.
D.1 Experimental Settings and Analysis
We recruited 30 volunteers with natural language interaction experience, each completing 10 sets of multi-domain dialogs. The total 300 dialog goals were selected from the final goals of successful dialogs in MultiWOZ, with a maximum of 40 turns consistent with the simulated experiment settings. The Table 5 presents the performance comparison of various methods in real user experiments.
DyBBT consistently achieved the highest task success rate and the lowest average dialog turns in real user experiments, corroborating the findings from simulated environments. This alignment demonstrates that the cognitive state space $\mathcal{C}$ effectively captures dynamic affordances in authentic dialog contexts. Furthermore, DyBBT attained the highest user satisfaction ratings, with qualitative feedback indicating that its dialog decisions were perceived as more natural and efficient. In contrast, variants employing random switching and other baseline methods exhibited significant deficiencies in both decision quality and response appropriateness.
Our analysis of the cognitive state distribution $\mathbf{c}_{t}=[d_{t},u_{t},\rho_{t}]$ in real dialogs revealed high consistency with the simulated environment distribution, validating the representativeness of our cognitive state assumptions in practical scenarios. The majority of failure cases were attributable to user intent shifts, irrelevant user responses, or non-standard linguistic input, challenges inherent to real world dialog systems that highlight areas for future improvement.
D.2 Case Studies
To provide qualitative insights into DyBBT’s performance with real users, we present three representative dialog cases from our experiments.
D.2.1 Case 1: Mid-Dialog User Intent Shift
The user initially requested a “Chinese food” restaurant reservation but suddenly asked “Are there any attraction nearby?” midway through the dialog.
Cognitive State:
  • $d_{t}=0.4$ (mid-stage)
  • $u_{t}=0.6$ (original goal incomplete; new intent introduces uncertainty)
  • $\rho_{t}=0.3$ (low dependency; cross domain intent)
DyBBT’s Decision:
  • System 1’s confidence $p_{t}^{S1}=0.55<\kappa$, triggering System 2.
  • System 2 generated multi-path strategies: 1.Confirm whether to continue the original task; 2.Provide attraction information; 3.Confirm to end the original task and ask attraction information.
Outcome: System 2 selected the “confirmation first” path, outputting ‘confirm(restaurant, food, “Chinese”)’ → ‘inform(attraction, type, “cinema”)’. The user confirmed continuing the original task, and the dialog was successfully completed.
Analysis: This case demonstrates DyBBT’s robustness in handling sudden user intent shifts. Although $u_{t}$ and $\rho_{t}$ in the cognitive state did not directly reflect the severity of the intent conflict, System 1’s low confidence accurately captured the uncertainty of the current strategy, thereby triggering System 2’s in-depth reasoning. Through multi-path evaluation, System 2 identified potential goal confusion and prioritized confirmation over directly responding to the new request, avoiding task interruption risks while maintaining dialog coherence via proactive clarification. This highlights the core advantage of the dual system architecture in addressing unexpected user behaviors.
D.2.2 Case 2: Vague User Information
When requesting a taxi, the user said “I want to go to that famous square” without specifying the exact name.
Cognitive State:
  • $d_{t}=0.2$ (early stage)
  • $u_{t}=0.9$ (high uncertainty)
  • $\rho_{t}=0.7$ (strong dependency between destination and departure location)
DyBBT’s Decision:
  • The cognitive state region had a low visitation count ($n_{t}(\mathbf{c}_{t})=8<\tau\sqrt{\log T}$), triggering System 2.
  • System 2 generated strategies: 1.Request the exact name; 2.List possible options; 3.Confirm the area.
Outcome: System 2 selected the “option list” path, outputting ‘request(taxi, destination)’ → ‘inform(attraction, name, “Central Square”)’. The user made a selection, and the task proceeded.
Analysis: This case highlights the value of the exploration condition in addressing vague user expressions. While simulated users typically provide explicit slot values, real world users often use vague references, which can easily stall standard strategies. DyBBT identified the unfamiliarity of this cognitive state through low visitation counts, activating System 2. The final option list strategy balanced information gaps and user experience, avoiding the poor experience caused by mechanical questioning while constraining the problem space through limited options. This proves that the exploration mechanism based on cognitive state visitation frequency can effectively identify dialog patterns not covered in simulated training and enhance the system’s adaptability in real scenarios via planned exploration.
D.2.3 Case 3: Non-Typical User Behavior
After completing a hotel reservation, the user suddenly repeatedly asked “Does the price include breakfast?”.
Cognitive State:
  • $d_{t}=0.9$ (late stage)
  • $u_{t}=0.1$ (low uncertainty; all slots confirmed)
  • $\rho_{t}=0.2$ (low dependency)
DyBBT’s Decision:
  • The cognitive state was judged as “well explored” with high System 1 confidence ($p_{t}^{S1}=0.88$), so System 2 was not triggered.
  • System 1 directly responded with ‘inform(hotel, breakfast, “no”)’.
Outcome: The user expressed dissatisfaction, perceiving the system’s response as “mechanical repetition”.
Analysis: This case reveals the limitations of the current cognitive state representation. The three dimensions cannot capture emotional factors behind users’ repeated questions. The system failed to recognize its unconventionality and the meta-controller missed the opportunity to trigger System 2, leading the system to respond in a standard but insufficiently empathetic manner. When user behaviors significantly deviate from the distribution of training data, the system lacks the ability to understand deeper semantic and emotional contexts in dialogs.
Appendix E Further Experimental Analysis
E.1 Experimental Results on MultiWOZ
Table 6 presents DyBBT’s performance on the MultiWOZ multi-domain dialog dataset, including key metrics (Inform, Success, Book, Turns). Compared with additional LLM based methods, it further validates DyBBT’s generalization ability and effectiveness.
[…]
Table 6: Evaluation results on MultiWOZ dataset. DyBBT-8B/GPT-4.0 denotes Qwen3-8B for System 1 and GPT-4.0 for System 2. DQN, LLM_DP and EIERL are reported in EIERLZhao et al. (2025), other results were reported from original papers, “—” indicates unreported results.
E.2 Ablation Study Settings and Results
This subsection details the settings of ablation studies and corresponding result tables, aiming to systematically validate the contributions of each core component of the DyBBT framework to overall performance. We conduct comprehensive ablation studies to evaluate the contribution of each component of the DyBBT framework on the MultiWOZ dataset, and the results are shown in Table 2:
  • DyBBT w/o MC: Replaces the meta-controller with random switching (each turn has a $10\%$ chance to invoke System 2).
  • DyBBT w/o S2: A degraded system that only uses System 1.
  • DyBBT w/o KD: Disables the knowledge distillation process. System 1 is never updated with data from System 2.
  • DyBBT w/o EC: Removes the exploration condition 1: ($n_{t}(\mathbf{c}_{t})<\tau\sqrt{\log T}$). System 2 is only triggered by low confidence (Condition 2).
  • DyBBT w/o CS: Replaces the cognitive state $\mathbf{c}_{t}$ with the raw, high dimensional belief state $\mathbf{s}_{t}$ (one-hot encoding of slot-values) for the meta-controller’s condition 1. The visitation count $n_{t}$ is computed over a discretized version of $\mathbf{s}_{t}$.
  • DyBBT w/o CC: Removes the confidence condition 2: ($p_{t}^{S1}<\kappa$). System 2 is only triggered by under explored states (Condition 1).
  • DyBBT w/ Learned CS: Replaces the hand-designed cognitive state $\mathbf{c}_{t}=[d_{t},u_{t},\rho_{t}]$ with a three dimensional embedding learned by a small MLP (2 layers, 32 units each) from the raw belief state $\mathbf{s}_{t}$. This tests the necessity of our specific cognitive state design.
  • DyBBT w/o $d_{t}$, w/o $u_{t}$, w/o $\rho_{t}$: Ablation studies removing one dimension from the cognitive state at a time to quantify its individual contribution.
E.3 Confidence Condition Error Analysis
To further clarify the crucial role of the Confidence Condition (CC) in the DyBBT framework, we conducted an in depth analysis of the types and proportions of errors prevented by this mechanism. The CC primarily serves as a safety net to prevent System 1 from making “catastrophic errors” in states with “high cognitive uncertainty”, whereas the absence of the Exploration Condition (EC) mainly leads to reduced “exploration efficiency” rather than direct task failures.
E.3.1 Types and Proportions of Errors Prevented by the CC
We analyzed a Sample of 200 CC interventions dialog logs of "DyBBT w/o EC" and "DyBBT w/o CC". Table 7 summarizes the distribution of error types among these cases.
Error Type Description Proportion Impact Level 1. Logical Conflict System 1’s proposed action contradicts the confirmed belief state 32% High 2. Context Mismatch System 1’s action is grammatically correct but inconsistent with the current dialog phase or user expectations 28% Low 3. Critical Information Omission System 1 fails to identify the next key slot necessary to complete the task 25% Medium 4. Domain/Slot Confusion System 1 confuses slots or selects the wrong domain in cross domain scenarios 15% High
Table 7: Types and proportions of errors prevented by the Confidence Condition
Types 1 and 4 account for 47% of errors, which are relatively severe and would almost certainly lead to dialog failure if not corrected by the CC. In contrast, System 2 invocations triggered by the EC are primarily used to explore unknown states to find better paths, and the cost of its “misses” is usually increased dialog turns rather than direct failure. This explains why removing the CC results in a more significant performance decline in ablation studies.
E.3.2 Case Analysis of CC Interventions
The following four cases demonstrate how the CC prevents serious errors in practice.
Case 1: CC Prevents a “Logical Conflict” Error
  • Background: After the user booked a restaurant, they requested a taxi.
  • Belief State:
  • System 1’s Incorrect Action: request(taxi, destination) with confidence $p_{t}^{S1}=0.62$ (low confidence, triggering CC)
  • CC’s Intervention: System 2 identified the logical flaw: the destination should be the restaurant, and the departure location should be queried instead. System 2 generated the correct action sequence: request(taxi, departure).
  • Analysis: System 1 made a fundamental logical error. Without CC correction, the system would have asked an unprocessable question, causing user confusion and most likely task failure. The CC acts as a commonsense reasoning validator here.
Case 2: CC Prevents a “Context Mismatch” Error
  • Background: Late in the dialog, all user constraints were satisfied, and the hotel was successfully booked.
  • Belief State:
  • System 1’s Suboptimal Action: request(hotel, parking) with confidence $p_{t}^{S1}=0.58$ (low confidence, triggering CC)
  • CC’s Intervention: System 2 determined that continuing to request information after task completion was redundant and confusing. The optimal action was to politely end the dialog. System 2 output: inform(booking, complete) $\rightarrow$ goodbye().
  • Analysis: System 1 failed to perceive the “conclusiveness affordance” of the dialog phase. The CC’s intervention prevented unnecessary interactions that could have reduced user experience, demonstrating its precise grasp of dialog rhythm and user expectations.
Case 3: CC Prevents a “Critical Information Omission” Error
  • Background: The user is booking a restaurant, having provided the number of people and date but not the time.
  • Belief State:
  • System 1’s Incorrect Action: inform(restaurant, name, "Pasta Place") with confidence $p_{t}^{S1}=0.55$ (low confidence, triggering CC)
  • CC’s Intervention: System 2 reasoned that the necessary booking time must be obtained before providing the restaurant name; otherwise, the booking process cannot be completed. System 2 generated the action sequence: request(restaurant, time).
  • Analysis: System 1 attempted to skip a critical step and directly provide information, which would have led to subsequent booking failure. The CC ensures information completeness on the critical path of the task, preventing interruptions to the task flow.
Case 4: CC Prevents a “Domain/Slot Confusion” Error
  • Background: A multi-domain scenario where the user wants to take a taxi from the hotel they are staying at.
  • Belief State:
  • System 1’s Incorrect Action: request(hotel, departure) with confidence $p_{t}^{S1}=0.60$ (low confidence, triggering CC)
  • CC’s Intervention: System 2 accurately identified that “departure” is a slot in the taxi domain, not an attribute of the hotel domain. System 2 corrected the action to: request(taxi, departure).
  • Analysis: System 1 confused slots across different domains, generating an invalid semantic action. Leveraging its stronger reasoning capabilities, the CC corrected this cross-domain understanding error, which is crucial in complex multi-turn, multi-domain dialogs.
In summary, the Confidence Condition is a crucial robustness safeguard mechanism in the DyBBT framework, which specifically targets the inherent weaknesses of System 1 when facing partial observability, logical conflicts, and context transitions. These errors are not only common but also fatal in nature. Hence, removing the CC causes a more severe performance decline than removing the EC in ablation experiments.
E.4 Supplementary Analysis Figures
This subsection provides all supplementary figures supporting the main text analysis in Sec. 5, which offer intuitive data support for the discussions:
  • Fig. 5: Heatmap of visitation frequency in the cognitive state space $\mathcal{C}$, illustrating the structured exploration strategy of the meta-controller across dialog phases.
  • Fig. 5: Analysis of meta-controller decisions, showing the rate of System 2 invocation across dialog progress and the proportion of triggers from each condition.
  • Fig. 7: Demonstrates the improvement of System 1 through knowledge distillation and the corresponding reduction in System 2 invocation over training.
  • Fig. 7: Compares the empirical cumulative regret of DyBBT against the theoretical upper bound derived under simplifying assumptions.
Figure 4: Visitation frequency in cognitive state space $\mathcal{C}$, showing the meta-controller’s phase-dependent exploration strategy across dialog progress and user uncertainty dimensions.
Figure 5: Analysis of meta-controller decisions. Rate of System 2 invocation across dialog progress. Pie chart showing the proportion of System 2 invocations.
Figure 6: System 1 improvement through knowledge distillation, which leads to monotonic improvement of System 1 and a corresponding reduction in the need to invoke System 2.
Figure 7: Empirical cumulative regret of DyBBT compared to the theoretical upper bound derived under simplifying assumptions. The sublinear growth of empirical regret is consistent with the theoretical intuition.
E.5 Hyperparameter Sensitivity Analysis
A key concern is the sensitivity of DyBBT’s performance to the meta-controller’s hyperparameters: the exploration threshold $\tau$, the confidence threshold $\kappa$, and the number of bins used to discretize the cognitive state space $\mathcal{C}$. We conducted a comprehensive grid search over $\tau\in\{0.5,1.0,1.5,2.0\}$, $\kappa\in\{0.5,0.6,0.7,0.8,0.9\}$, and bin counts $\in\{3,4,5,6,7\}$ on both the MS Dialog and MultiWOZ development sets. Performance is measured by the success rate (%), and the results are visualized in Fig. 8.
Figure 8: 3D surface plots of success rate (%) as a function of $\tau$, $\kappa$, and bin count for (left) MS Dialog and (right) MultiWOZ. The optimal configuration ($\tau=1.0$, $\kappa=0.7$, $bins=5$) is marked with a red star.
The results indicate that DyBBT is robust to a wide range of hyperparameter choices. High performance (success rate $83\%$ in MS Dialog and $82\%$ in MultiWOZ) is sustained within the region $\tau\in[0.8,1.2]$, $\kappa\in[0.6,0.8]$ and bin count $\in[4,6]$. The chosen values ($\tau=1.0$, $\kappa=0.7$, $bins=5$) lie at the center of this high performance plateau, achieving $86.1\%$ average on MS Dialog and $84.1\%$ on MultiWOZ. This configuration maximizes both performance and robustness across domains.
We also observe that the bin count has a moderate impact on performance. Too few bins oversimplify the cognitive state, leading to under exploration; too many bins increase the risk of overfitting and reduce the effectiveness of the visitation count. A bin count of 5 strikes an optimal balance, capturing sufficient state granularity without sacrificing generalization.
E.6 Model Scaling Analysis
Model Family Size Params Success Rate $\uparrow$ Inference Time $\downarrow$ Cost-Effectiveness $\uparrow$ Llama-3.2 1B 1.1B $78.3\pm 0.017$ 0.32x 244.7 3B 3.0B $80.1\pm 0.015$ 0.48x 166.9 7B 6.7B $81.9\pm 0.013$ 0.75x 109.2 8B 8.0B $82.6\pm 0.012$ 0.89x 92.8 Qwen2.5 0.5B 0.5B $77.4\pm 0.018$ 0.28x 276.4 1.5B 1.7B $79.6\pm 0.016$ 0.41x 194.1 3B 2.9B $81.5\pm 0.014$ 0.59x 138.1 7B 6.6B $83.1\pm 0.012$ 0.86x 96.6 Qwen3 0.6B 0.6B $79.2\pm 0.016$ 0.35x 226.3 1.7B 1.8B $81.2\pm 0.014$ 0.52x 156.1 4B 4.2B $83.6\pm 0.011$ 0.78x 107.2 8B 8.0B $\mathbf{85.1\pm 0.011}$ 1.00x 85.10
Table 8: Model scaling analysis across three model families on MultiWOZ 2.1. Success Rate is reported with standard deviation over 5 seeds. Inference Time is normalized to Qwen3-8B (1.0x)
To systematically evaluate the impact of model scale on DyBBT’s performance and efficiency, we conduct a comprehensive scaling analysis using three prominent open weight model families: Llama-3.2 Instruct(1B-8B), Qwen2.5 Instruct(0.5B-7B), and Qwen3 (0.6B-8B) on the MultiWOZ 2.1 benchmark. Performance is measured by Success Rate and Inference Time relative to Qwen3-8B, Cost-Effectiveness is defined as Success Rate divided by Inference Time. Results are summarized in Table 8.
The results reveal several key trends. First, across all model families, larger models consistently achieve higher success rates, demonstrating the benefit of increased capacity for both intuitive response generation (System 1) and deliberative reasoning (System 2). Second, at similar parameter scales, Qwen3 models outperform their Qwen2.5 counterparts, which in turn outperform Llama-3.2 models. This hierarchy aligns with the established capabilities of these families on reasoning intensive tasks.
These performance gains come with increased computational cost. Qwen3 models exhibit the longest inference times due to their architectural optimizations for complex reasoning, a cost further amplified when System 2 activates the model’s internal “think” mode for deliberate planning. Consequently, while Qwen3-8B delivers the highest absolute performance, its cost effectiveness (0.851) is lower than that of smaller models. Among the larger models, Qwen2.5-7B offers a favorable balance, achieving 97.6% of the performance of Qwen3-8B at 86% of the inference cost.
This analysis underscores a critical trade-off in deploying DyBBT: model scale must be chosen based on the specific application’s requirements for both performance and latency. For high stakes scenarios demanding maximum success rates, Qwen3-8B is the superior choice. For applications where computational efficiency is prioritized, a medium scale model like Qwen2.5-7B or Qwen3-4B provides a highly competitive performance cost ratio.
E.7 Cost Effectiveness Analysis of Different System Configurations
To provide practitioners with a clear cost performance trade off analysis, we compare DyBBT-8B, DyBBT-8B/GPT-4.0, and LLM_DP (pure GPT-4.0) on the MultiWOZ dataset. Since GPT-4.0 is only available via commercial APIs, we adopt two alternative evaluation approaches: measuring end-to-end inference time under the same hardware environment, and calculating economic cost based on actual token usage.
All local models run on an NVIDIA 5090 GPU, while the API model (GPT-4.0) is accessed via the official interface. The end-to-end Inference Time including model forward propagation or API call latency, averaged seconds per dialog. Normalized Inference Time is benchmarked against DyBBT-8B’s inference time. API Cost is based on GPT-4.0’s official pricing (input: \$0.03 per 1k tokens; output: \$0.06 per 1k tokens).
[…]
Table 9: Cost effectiveness analysis of different system configurations
Table 9 presents the comprehensive cost effectiveness comparison. Compared to DyBBT-8B, DyBBT-8B/GPT-4.0 achieves only a 1.2% improvement in success rate, but incurs a 2.3× increase in inference time and a cost of \$0.16 per dialog. This indicates that marginal performance gains are accompanied by substantial computational overhead and economic costs. LLM_DP (GPT-4.0), which relies solely on well designed prompts to enable LLMs to generate system actions, not only achieves an extremely low success rate but also has the longest inference time and highest API cost, highlighting the advantage of the DyBBT framework in balancing performance and cost. The System 2 invocation ratio of DyBBT-8B/GPT-4.0 is only 14.3%, indicating that the Meta-Controller effectively limits the use of expensive APIs. However, API call latency still dominates the total inference time.
In practical deployment scenarios, if ultimate performance is pursued and API dependency/latency is acceptable, using GPT-4.0 or more advanced closed source models for System 2 is an option. This requires balancing the 1.2% performance gain against the 2.3× inference time and additional costs. Since DyBBT already achieves excellent performance at the 8B scale, DyBBT-8B offers the optimal trade-off when computational efficiency, independence, and cost effectiveness are prioritized.
E.8 Comparison with Qwen3’s Native Switching
Configuration Success Rate $\uparrow$ Normalized Time $\downarrow$ Cost Effectiveness $\uparrow$ S1 no think / S2 no think $79.6\pm 0.015$ 0.6x 132.7 S1 think / S2 think $\mathbf{86.5\pm 0.010}$ 3.2x 27.0 DyBBT (S1 no think / S2 think) $85.1\pm 0.011$ 1.0x 85.1
Table 10: Comparison between DyBBT’s meta-controller and Qwen3’s native switching mechanism. Normalized time is normalized to DyBBT’s default mode (S1 no think / S2 think = 1.0x).
To further validate the effectiveness of DyBBT’s bandit inspired meta-controller, we compare it against the native fast/think mode switching mechanism built into Qwen3-8B. Qwen3 natively supports a heuristic switching logic based on its internal confidence estimation, allowing it to dynamically activate a more expensive “think” mode for complex reasoning. We evaluate three configurations:
  1. 1.
    S1 no think / S2 no think: Both systems use the standard forward pass without activating Qwen3’s internal think mode.
  2. 2.
    S1 think / S2 think: Both systems always use the think mode, representing a high cost, high deliberation baseline.
  3. 3.
    S1 no think / S2 think: DyBBT’s mode, System 1 operates in fast mode, while System 2 uses think mode when triggered by the meta-controller.
We report performance on the MultiWOZ test set also using Success Rate, Inference Time (with DyBBT’s default mode as 1.0x), and Cost-Effectiveness Results are summarized in Table 10.
As anticipated, the always think configuration achieves the highest success rate (86.5%), confirming that maximal deliberation improves task performance. However, this comes at an prohibitive computational cost 3.2× the inference time of the selective activation of DyBBT. In contrast, DyBBT’s mode achieves nearly comparable performance (85.1% success) with only one-third of computational overhead, resulting in a significantly higher cost-effectiveness.
The no-think baseline performs poorly, underscoring the necessity of deliberate reasoning in complex dialog states. DyBBT strikes a balance between these extremes by invoking costly reasoning only when cognitively justified, either due to under exploration or low confidence, leading to near optimal performance with moderate and targeted computational overhead. This leads to less efficient allocation of computational resources, as also reflected in human evaluation (Sec. 4.4).
E.9 Failure Mode Analysis and Limitations
While DyBBT demonstrates strong performance across benchmarks, we conducted a comprehensive failure mode analysis to understand its limitations in practical deployment scenarios. Through post-hoc analysis on 1000 dialogs of MultiWOZ with cross validation by three expert annotators, we quantitatively assessed the occurrence rates of different failure modes.
Category Description Rate Impact Level
Inaccurate Cognitive State Representation Handcrafted $\mathbf{c}_{t}$ fails to capture complex dialog dynamics like abrupt intent shifts 3.1% High
Propagation of System 2 Demonstration Errors Errors in System 2’s reasoning or self evaluation distilled into System 1 1.4% Medium
Underexploration Due to State Discretization Heuristic quantization of $\mathcal{C}$ masks critical state differences 0.7% Low
Total Failure Rate 5.2%
Table 11: Quantitative analysis of DyBBT failure modes on MultiWOZ dataset (N=1000 dialogs)
Table 11 presents the quantitative breakdown of failure modes, revealing that 94.8% of dialogs proceed without significant failures while only 0.3% exhibit multiple concurrent failure modes. The failure modes primarily occur in edge cases characterized by abrupt user intent shifts, complex cross domain dependencies, and non-standard user behaviors. These scenarios constitute inherently challenging “hard cases” that represent a minority in real world task oriented dialogs. The built-in safety mechanisms demonstrate substantial protective value: the Confidence Condition intercepts 76% of System 1’s low confidence errors, preventing catastrophic failures in uncertain states; Knowledge Distillation reduces System 2 invocation rate by 42% (Fig. 7), progressively mitigating error propagation risks; and human evaluation shows 88.7% alignment with expert judgment, far exceeding the random switching baseline. These builtin safety mechanisms demonstrate substantial protective value.
For the majority of commercial task oriented dialog scenarios, DyBBT’s current failure profile represents an acceptable risk given its significant performance advantages. However, in safety critical domains, the identified failure modes warrant additional safeguards. Our future work addresses these limitations through end-to-end learned cognitive representations, improved uncertainty calibration, and adaptive exploration mechanisms. These evolutionary improvements will further enhance DyBBT’s robustness while preserving its core architectural advantages for practical deployment.
E.10 Case Study
To qualitatively validate the efficacy of the meta-controller’s switching mechanism beyond aggregate metrics, we present contrasting case studies sampled from the MultiWOZ test set. These examples illustrate how DyBBT’s principled switching aligns with human judgment in successful cases, and reveal its limitations in failure scenarios, providing concrete insights into the operational boundaries of our framework.
E.10.1 Case 1: Successful Intervention due to High Epistemic Uncertainty
This case demonstrates the meta-controller correctly triggering System 2 for targeted exploration in a novel cognitive state, leading to successful task completion.
Belief State Context:
Cognitive State Analysis:
  • Dialog Progress ($d_{t}$): 0.15 (Early stage, 6/40 turns)
  • User Uncertainty ($u_{t}$): 0.8 (High, 4 out of 5 key slots unconfirmed or unknown)
  • Slot Dependency ($\rho_{t}$): 0.6 (Medium, ‘area’ and ‘name’ often co-occur in restaurant domain)
Meta-Controller Decision: The visitation count for this cognitive state region was low ($n_{t}(\mathbf{c}_{t})=12<\tau\sqrt{\log T}\approx 25$), triggering System 2 via the exploration condition. System 1’s confidence was medium ($p_{t}^{S1}=0.75>\kappa$). System 2 Intervention: System 2 performed multi-path reasoning. The top ranked sequence prioritized gathering the uncertain location information: request(restaurant, area)confirm(restaurant, area, “north”)inform(restaurant, name, “Golden Dragon”). Outcome: This strategy efficiently disambiguated the user’s intent. The dialog was successfully completed 6 turns later. This case exemplifies how DyBBT’s exploration condition actively targets under explored regions of $\mathcal{C}$ for strategic information gain, a key advantage over static exploration policies.
E.10.2 Case 2: Successful Intervention due to Low Aleatoric Confidence
This case highlights the robustness safeguard of the confidence condition, preventing a potential failure due to System 1’s overconfidence in a complex state.
Belief State Context:
Cognitive State Analysis:
  • Dialog Progress ($d_{t}$): 0.65 (Late stage, 26/40 turns)
  • User Uncertainty ($u_{t}$): 0.1 (Low, most slots confirmed)
  • Slot Dependency ($\rho_{t}$): 0.9 (High, ‘train’ and ‘taxi’ domains are logically connected)
Meta-Controller Decision: The cognitive state was well explored ($n_{t}(\mathbf{c}_{t})=505$). However, System 1’s confidence was low ($p_{t}^{S1}=0.55<\kappa=0.7$), likely due to the logical conflict between the train destination (“cambridge”) and the taxi destination (“london stansted airport”). This triggered System 2 via the confidence condition. System 2 Intervention: System 2 identified the inconsistency. Its reasoning path prioritized clarification: confirm(train, destination, “cambridge”)confirm(taxi, destination, “london stansted airport”)request(clarification, “final destination”). Outcome: The user clarified they meant “Cambridge” was the departure point for the taxi, not the destination. This intervention corrected a critical misunderstanding that would have led to task failure. This case underscores the critical role of the confidence condition in mitigating System 1’s limitations and handling partial observability.
E.10.3 Case 3: Failure due to Cognitive State Misrepresentation
This case illustrates a fundamental limitation: the handcrafted cognitive state can fail to capture critical dialog nuances, leading to a suboptimal decision.
Belief State Context:
Cognitive State Analysis:
  • Dialog Progress ($d_{t}$): 0.8 (Late stage, booking complete)
  • User Uncertainty ($u_{t}$): 0.4 (Medium, ‘attraction/name’ unknown)
  • Slot Dependency ($\rho_{t}$): 0.7 (High, ‘hotel/area’ and ‘attraction/area’ match)
Meta-Controller Decision: The state had medium visitation ($n_{t}(\mathbf{c}_{t})=162$) and System 1 was highly confident ($p_{t}^{S1}=0.92$) in its action to request(attraction, name). The meta-controller did not trigger System 2. Analysis of Failure: While the cognitive state suggested a routine information gathering context, it failed to capture the user had just finished a complex booking and was likely expecting a concise recommendation, not another request. The best policy should afford an inform(attraction, name, “museum of science”) action. Outcome: This case reveals the limitation of fixed, hand engineered cognitive features and points to the need for more adaptive or learned state representations in future work.
E.10.4 Summary and Limitations
These case studies provide concrete evidence that DyBBT’s meta-controller dynamically allocates computational resources in a manner that is both effective and efficient, closely mirroring human expert judgment in successful cases (Cases 1 & 2). The failures (Case 3) are highly instructive, revealing that the primary limitation lies not in the switching mechanism itself, but in the fidelity of the handcrafted cognitive state $\mathbf{c}_{t}$ to represent all critical aspects of the dialog context. Future work will focus on learning this state representation end-to-end from data, which could mitigate such representational gaps and further enhance the framework’s robustness and applicability.